If your p-value is 0.33 and your alpha level is 0.1 then you would:

  • reject the Ho
  • fail to reject the Ho
  • unable to determine from this information
  • accept the Ho



If your p-value is 0.33 and your alpha level is 0.05 then you would:

  • reject the Ho
  • fail to reject the Ho
  • unable to determine from this information
  • accept the Ho



If your p-value is 0.33 and your alpha level is 0.01 then you would:

  • reject the Ho
  • fail to reject the Ho
  • unable to determine from this information
  • accept the Ho



If your p-value is 0.33 and your alpha level is 0.005 then you would:

  • reject the Ho
  • fail to reject the Ho
  • unable to determine from this information
  • accept the Ho



If your p-value is 0.33 and your alpha level is 0.001 then you would:

  • reject the Ho
  • fail to reject the Ho
  • unable to determine from this information
  • accept the Ho



If your p-value is 0.33 and your alpha level is 0.0005 then you would:

  • reject the Ho
  • fail to reject the Ho
  • unable to determine from this information
  • accept the Ho



If your p-value is 0.33 and your alpha level is 0.0001 then you would:

  • reject the Ho
  • fail to reject the Ho
  • unable to determine from this information
  • accept the Ho



If your p-value is 0.33, what is the most significant alpha level?

  • There is no 'most significant alpha level'
  • 0.1
  • 0.05
  • 0.01
  • 0.005
  • 0.001
  • 0.0005
  • 0.0001
Answer #1

Answer

when p value is less than the alpha level , then rejecting the null hypothesis (H0), if p value is greater than alpha , then do not rejecting H0.

1) p value = 0.33

\alpha = 0.1

since p value = 0.33 > \alpha = 0.1

so , fail to reject H0

2)

p value = 0.33

\alpha = 0.05

since p value = 0.33 > \alpha = 0.05

so , fail to reject H0

3)

p value = 0.33

\alpha = 0.01

since p value = 0.33 > \alpha = 0.01

so , fail to reject H0

4)

p value = 0.33

\alpha = 0.005

since p value = 0.33 > \alpha = 0.005

so , fail to reject H0

5)

p value = 0.33

\alpha = 0.001

since p value = 0.33 > \alpha = 0.001

so , fail to reject H0

6)

p value = 0.33

\alpha = 0.0005

since p value = 0.33 > \alpha = 0.0005

so , fail to reject H0

7)

p value = 0.33

\alpha = 0.0001

since p value = 0.33 > \alpha = 0.0001

so , fail to reject H0

8)

significance alpha level is the probability of rejecting th null hypothesis whn it is true,]

since given p value is greater than the all provided alpha level so,

there is no "most significant alpha level".